# Amputation-imputation based generation of synthetic tabular data for ratemaking

### v1

@@ Abstract
Actuarial ratemaking depends on high-quality data, yet access to such data is often limited by the cost of obtaining new data, privacy concerns, etc. In this paper, we explore synthetic-data generation as a potential solution to these issues. In addition to discussing generative methods previously studied in the actuarial literature, we introduce to the insurance community another approach based on Multiple Imputation by Chained Equations (MICE). We present a comparative study using an open-source dataset and evaluating MICE-based models against other generative models like Variational Autoencoders and Conditional Tabular Generative Adversarial Networks. We assess how well synthetic data preserves the original marginal distributions of variables as well as the multivariate relationships among covariates. We also investigate the consistency between Generalized Linear Models (GLMs) trained on synthetic data with GLMs trained on the original data. Furthermore, we assess the ease of use of each generative approach and study the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim counts. Our results highlight the potential of MICE-based methods in creating high-quality tabular data while being more user-friendly than the other methods.
Keywords: synthetic actuarial data, data augmentation, multiple imputation by chained equations, generative adversarial networks, autoencoders
@@ 1 Introduction
Motivation. High-quality actuarial data is essential for actuaries in various business lines of insurance and reinsurance companies. It helps quantify risks more accurately, and thus maintain the competitiveness of actuarial pricing and reserving models. Furthermore, researchers in actuarial science need realistic data to test their novel methodologies and compare them with existing ones. Even if insurers and reinsurers may have enough data for some products or business lines, the companies rarely publicly share their data due to concerns related to data privacy, security, and their competitive advantage. As a result, the number of open-access actuarial datasets is quite small (see, e.g., Gabrielli and Wüthrich, (2018), Cote et al., (2020), Segura-Gisbert et al., (2024)). Furthermore, the companies themselves may also face a lack of data, e.g., when they enter a new market or release a new (re-)insurance product.
@@
To increase the amount of actuarial data, one can generate synthetic data, which can be used as a standalone dataset or as an enlargement to the existing dataset. In the former case, synthetic data can be publicly shared and used for research purposes (e.g., testing new models and algorithms), if it has the statistical properties of the real data and does not expose sensitive information. As for the latter case, the augmentation of real data by synthetic data allows to train models on a larger and potentially more diverse dataset and, as a result, improve their capacity to generalize to unseen data. It can be especially helpful for actuaries when collecting and processing additional real data is impossible or very costly in terms of time or money.
@@
There are two types of approaches to generating synthetic data: those that have assumptions regarding the distribution of data to be generated and those that do not impose such assumptions. We will focus on the latter type, as one of our aims is to develop an actuarial data generator that can be easily used “out-of-the-box”. Furthermore, the insurance application we target is general ratemaking (without telematics), where, usually, datasets are tabular with both numeric and categorical variables and temporal independence of observations is assumed.
@@
The generation of synthetic data without the need for distributional assumptions gained much popularity due to the advancement of deep generative models such as generative adversarial networks (GANs) and variational autoencoders (VAEs). To our knowledge, only three papers investigated the use of such models in the context of ratemaking: Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024). We discuss these articles in more detail in the literature overview below and mention for now only aspects that strongly motivate us to conduct further research on this topic. First, in our experience, the data generation approaches studied in those three papers seem to require quite some customization for each input dataset, which is why their ease of use in daily work of pricing actuaries seems to be limited. Second, the benefit of augmenting original data with synthetic data is not addressed in those papers.
Our contributions. First, we introduce to actuarial science several amputation-imputation-based approaches to data generation. These techniques are based on Multiple Imputation by Chained-Equations (MICE) with Random Forests (RFs) as imputation models. According to our tests, these approaches can be more suitable for “out-of-the-box” usage in practice than the generative models studied in Kuo, (2020), Cote et al., (2020), Jamotton and Hainaut, (2024), while having a comparable performance.
Second, using a publicly available French Motor Third-Part Liability dataset called freMTPL2freq (Dutang and Charpentier,, 2019) as a training dataset, we compare $9$ approaches to ratemaking data generation, including the existing approaches (Kuo,, 2020; Cote et al.,, 2020; Jamotton and Hainaut,, 2024), their modifications that address the challenge of generative adversarial networks (GANs) in dealing with high-cardinality categorical variables, and the newly introduced amputation-imputation-based methods. We analyze the quality of synthetic datasets by studying two aspects: the similarity between the distribution of the original data and of synthetic data, the ability of methods to capture structural dependence among covariates and the response variable. Furthermore, we assess the impact of augmenting original data with synthetic data on the performance of GLMs for predicting claim frequency.
Our main results are threefold. First, we illustrate that the MICE-RF approach is a competitive method for generation of synthetic ratemaking data in comparison to deep generative models such as a VAE and a CTGAN. The important practical advantage of the MICE-RF algorithm is its ease of use, given the existing packages in $\mathrm{R}$ and the versatility of RFs in dealing with variables of different types without much pre-processing. Second, we show that modifying a CTGAN with an AE for categorical-variable pre-processing improves the quality of synthetic high-cardinality categorical variables but deteriorates the performance of this hybrid approach in terms of other criteria. In general, a MICE-RF approach and a CTGAN without AE seem to be the most balanced in terms of the performance metrics we consider. Third, our case studies indicate that for our experimental data sets based on freMTPL2freq data augmentation does not improve the performance of GLMs trained on the original data merged with synthetic one.
@@
Literature overview. Kuo, (2020) pioneered the use of conditional tabular GANs for generating actuarial datasets, focusing on ratemaking and lapse modeling. In his experiments with freMTPL2freq dataset, the researcher showed that generalized linear models (GLMs) trained on synthetic data achieved performance comparable to GLMs trained on real data, and the distributions of both continuous and categorical variables in the synthetic datasets closely mirrored the real datasets. However, the relativities (how much each level of a categorical variable affects the predicted outcome compared to a reference level) of GLM models constructed using synthetic datasets do not show the same patterns as those of models created using real datasets. Because of the observed instability in the parameters of a GLM trained on synthetic data, developing a rating plan based on it might not be the best approach.
@@
Inspired by Kuo, (2020), Cote et al., (2020) investigated three GAN-based models for generating synthetic insurance data: Multi-Categorical Wasserstein GAN with Gradient Penalty (MC-WGAN-GP), Conditional Tabular GAN (CTGAN), and Mixed Numerical and Categorical Differentially Private GAN (MNCDP-GAN). In a case study using the freMTPLfreq data, they find that MC-WGAN-GP performs best in replicating univariate distributions and the observed claim frequency for each category of categorical variables, followed by CTGAN, and then MNCDP-GAN. In assessing the consistency of GLMs fitted on real versus synthetic data, CTGAN-based GLMs most closely resemble the real-data-based GLM in terms of average estimated GLM coefficients. However, MC-WGAN-GP based GLM predictions are closest to the real-data-based GLM predictions when using Mean Absolute Error and Mean Squared Error as evaluation metrics. The authors conclude that neither of the tested models performs well across all key performance indicators (KPIs) with a high level of training data confidentiality.
@@
As an alternative to GAN-based models, Jamotton and Hainaut, (2024) propose a Variational Autoencoder (VAE) model for the generation of insurance data. Their approach features a quantile transformation for continuous variables to address multimodality. In their VAE, the researchers use a reconstruction loss function that consists of two parts: cross-entropy for categorical variables and MSE for continuous variables. In a case study using a data set from the homepage of the book Ohlsson and Johansson, (2010), the researchers find that the marginal distributions of the real datasets and those of the synthetic dataset are almost the same, as supported by visual inspection, Kolmogorov-Smirnov tests for continuous variables, and Chi-Squared tests for categorical variables. Based on Pearson and Spearman correlations for continuous variables and contingency scores based on contingency tables for categorical variables, the authors conclude that synthetic data preserves pairwise relationships among variables reasonably well. Training a GLM on synthetic data generated by their VAE and noticing that the respective Poisson deviance is almost the same as the one for a GLM trained on the original data, the scientists conclude that synthetic data can substitute real data in actuarial analyses.
Fundamentally different data generation methods are those based on data imputation and have not yet been explored in the actuarial science context. The main idea behind those methods is to sequentially “amputate” parts of data and then impute it with a suitable statistical learning model. To our knowledge, this idea was first introduced by D.B. Rubin, who first developed a data imputation approach called MICE in the 1970s (cf. Rubin, (1977)) then proposed to use this approach also to create (fully or partially) synthetic datasets in Rubin, (1993). We refer readers interested in the subsequent development of MICE-based generation of synthetic data to Drechsler, (2011) and emphasize two aspects. First, this approach gained popularity in various scientific domains (e.g., see Volker and Vink, (2021) for medicine), yet it seems to have been overseen by the actuarial community. Second, the imputation in this framework is done on a variable-by-variable (column-by-column) basis, and the imputation model can be flexibly chosen from linear regression to sophisticated models.
@@
A related idea of generating artificial data using imputation appeared in Neves et al., (2022), where the authors propose a so-called tabulator framework. In their approach, they propose to randomly (not on variable-by-variable basis) select data entries to be amputated and then imputed by a so-called back-end engine for imputation. As a model driving this imputation engine the authors propose to use one of the three methods that can be seen as modifications of Generative Adversarial Imputation Networks (GAINs, originally proposed by Yoon et al., (2018)): Slim GAIN (SGAIN), Wasserstein Slim GAIN with Clipping Penalty (WSGAIN-CP), Wasserstein Slim GAIN with Gradient Penalty (WSGAIN-GP). According to the researchers, tabulator has a comparable performance (tested on 10 datasets from various fields like healthcare, economics, and computer science) with CTGAN that, in turn, is better than medGAN, VEEGAN, and table-GAN. The researchers conclude that the tabulator preserves the dependence structure of the original dataset better than CTGAN, but both approaches struggle when generating synthetic data from original imbalanced data. It is also stated that synthetic data from tabulator improves the performance of regression/classification tasks only marginally.
@@
Readers interested in the literature on data generation beyond traditional ratemaking are referred to So et al., (2021) and So and Jeong, (2022) for telematics data, to Gabrielli and Wüthrich, (2018), Avanzi et al., (2021), Avanzi et al., (2023), Bauwelinckx et al., (2024) for reserving data, to Campo and Antonio, (2023) for auto-insurance fraud-detection data, and to Chapter 10 in Wuthrich et al., (2025) for a brief introduction to generative modeling tools for actuaries. We refer readers to Fonseca and Bacao, (2023) and Brophy et al., (2023) for a systematic review of the literature on the GAN-based generation of tabular data and on the GAN-based generation of time series data, respectively. For an overview comparing deep generative models, including but not limited to VAEs and GANs, we direct readers to Bond-Taylor et al., (2022).
@@ 2 General setup
We denote a tabular dataset by $\mathcal{S}:=\left\{\bm{s}_{i}\right\}_{i=1,\dots,n_{ROD}}$, where $n_{ROD}$ is the number of rows in the original data (number of observations), $\bm{s}_{i}=(\bm{x}_{i},v_{i},y_{i})$ is a tuple with three elements such that $\bm{x}_{i}=(x_{i;1},\dots,x_{i;j},\dots,x_{i;n_{C}})^{\top}$ is a $n_{C}$-dimensional vector of observed values of covariates (independent variables), $n_{C}\in\mathbb{N}$, is the transpose operator, $v_{i}\in(0,1]$ is an exposure-to-risk (proportion of a year during which the observation $i$ was tracked) and $y_{i}$ is an observed value of an one-dimensional response variable (dependent variable) for anobservation $i=1,\dots,n_{ROD}$. A tuple $\bm{s}_{i}$ can be interpreted as an insurance policy snippet, $\bm{x}_{i}$ – as a vector of policy characteristics, $v_{i}$ – as the fraction of a year during which a policy snippet $i$ was active, and $y_{i}$ – as the observed number of claims (or the monetary loss associated with a claim in the case of severity modeling). For simplicity, we may omit the index $i$ indicating the ordinal number of the observation, when it is not essential for interpretation, and just use $(\bm{x},v,y)$. The vector $\bm{x}$ can have both categorical and numerical components. We denote by $n_{CC}$ the number of categorical covariates in $\bm{x}$ and by $n_{NC}=n_{C}-n_{CC}$ the number of numerical covariates in $\bm{x}$.
@@
In ratemaking, it is common to model the link between $y$ and $(\bm{x},v)$ via a generalized linear model (GLM). GLMs were introduced in Nelder and Wedderburn, (1972) and became very popular in insurance, where they are routinely used in various business lines such as claim management, underwriting, etc. In our paper, we will only discuss the basics of a Poisson GLM, which is frequently used to predict insurance claim counts and which we use in numerical studies. For a deep dive into the theory and practice of GLMs in insurance readers are referred to Chapter 4 of Denuit and Trufin, (2019) and Chapter 5 of Wüthrich and Merz, (2023).
@@
In (2.3), $x_{j}$ (and/or $x_{k}$) can be just a feature representing one category. To consider an interaction with an entire categorical variable, a sum of different (2.3) (one for each category in the respective categorical variable, except for the dummy one) must be added to the right-hand side of (2.2). For more information on modeling interactions in GLMs, see, e.g., Fox and Weisberg, (2018), or Section 4.2.5 in Denuit and Trufin, (2019).
Let us denote by $\bm{\beta}^{\ast}$ the true value of $\bm{\beta}$. In practice, $\bm{\beta}^{\ast}$ is not known and must be estimated from collected data. Using a training data $\mathcal{S}_{train}$, an actuary can obtain an estimate of $\bm{\beta}^{\ast}$ by a so-called maximum likelihood estimation (MLE), which is equivalent to minimizing Poisson deviance in the case of a Poisson GLM. When comparing different GLM models, the one with a lower Poisson deviance on the respective test set $\mathcal{S}_{test}:=\mathcal{S}\setminus\mathcal{S}_{train}$ is considered a better-performing one. An alternative performance measure is the Akaike Information Criterion (AIC), which takes into account the number of parameters of a GLM and is based only on $\mathcal{S}_{train}$.
@@
To address the aspect of data quantity and quality, an actuary can generate a synthetic dataset $\mathcal{S}^{A}$ of a desirable size and reasonable quality. In this situation, two natural questions arise: one is related to the usage of synthetic data as a standalone set, and the other one is related to working with the original dataset that is augmented with a part of the synthetic dataset. First, how well does $\mathcal{S}^{A}$ preserve (from training data $\mathcal{S}_{train}$) the marginal distributions of variables as well as the multivariate relationships among covariates and how consistent is a GLM trained on $\mathcal{S}^{A}$ with a GLM fitted on $\mathcal{S}_{train}$? Second, which advantages and disadvantages does training GLMs on an augmented dataset $\mathcal{S}_{train}\cup\mathcal{S}^{A}$ have in terms of GLM’s predictive performance and the estimates of $\bm{\beta}^{\ast}$ (e.g., $\widehat{\bm{\beta}}(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ and/or $\widehat{\widehat{\bm{\beta}}\,}\!(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ are closer to the true $\bm{\beta}^{\ast}$)? These are the two main research questions that we aim to answer in this paper.

### v2

# Synthetic data for ratemaking: imputation-based methods vs adversarial networks and autoencoders

@@ Abstract
Actuarial ratemaking depends on high-quality data, yet access to such data is often limited by the cost of obtaining new data, privacy concerns, etc. In this paper, we explore synthetic-data generation as a potential solution to these issues. In addition to generative methods previously studied in the actuarial literature, we explore and benchmark another class of approaches based on Multivariate Imputation by Chained Equations (MICE). In a comparative study using an open-source dataset, MICE-based models are evaluated against other generative models like Variational Autoencoders and Conditional Tabular Generative Adversarial Networks. We assess how well synthetic data preserves the original marginal distributions of variables as well as the multivariate relationships among covariates. The consistency between Generalized Linear Models (GLMs) trained on synthetic data with GLMs trained on the original data is also investigated. Furthermore, we assess the ease of use of each generative approach and study the impact of generically augmenting original data with synthetic data on the performance of GLMs for predicting claim counts. Our results highlight the potential of MICE-based methods in creating high-fidelity tabular data while offering lower implementation complexity compared to deep generative models.
Keywords: synthetic actuarial data, data augmentation, multivariate imputation by chained equations, generative adversarial networks, autoencoders
@@ 1 Introduction
Motivation. High-quality actuarial data is essential for actuaries in various business lines of insurance and reinsurance companies. It helps quantify risks more accurately, and thus maintain the competitiveness of actuarial pricing and reserving models. Furthermore, researchers in actuarial science need realistic data to test their novel methodologies and compare them with existing ones. Even if insurers and reinsurers may have enough data for some products or business lines, the companies rarely publicly share their data due to concerns related to data privacy, security, and their competitive advantage. As a result, the number of open-access actuarial datasets is quite small (see, e.g., Gabrielli and Wüthrich (2018), Côté et al. (2025), Segura-Gisbert et al. (2024)). Furthermore, the companies themselves may also face a lack of data, e.g., when they enter a new market or release a new (re-)insurance product.
@@
To increase the amount of actuarial data, one can generate synthetic data, which can be used as a standalone dataset or as an enlargement to an existing dataset. In the former case, synthetic data can be publicly shared and used for research purposes (e.g., testing new models and algorithms), if it has the statistical properties of the real data and does not expose sensitive information. As for the latter case, the augmentation of real data by synthetic data allows to train models on a larger and potentially more diverse dataset and, as a result, improve their capacity to generalize to unseen data. It can be especially helpful for actuaries when collecting and processing additional real data is impossible or very costly in terms of time or money.
@@
There are two types of approaches to generating synthetic data: those that have assumptions regarding the distribution of data to be generated and those that do not impose such assumptions. We will focus on the latter type, as one of our aims is to develop an actuarial data generator that can be easily used “out-of-the-box”. Furthermore, the insurance application we target is general ratemaking, where, usually, datasets are tabular with both numeric and categorical variables and temporal independence of observations is assumed.
@@
The generation of synthetic data without the need for distributional assumptions gained much popularity due to the advancement of deep generative models such as generative adversarial networks (GANs) and variational autoencoders (VAEs). To our knowledge, only three papers investigated the use of such models in the context of ratemaking: Kuo (2020), Côté et al. (2025), Jamotton and Hainaut (2024). We discuss these articles in more detail in the literature overview below and mention for now only aspects that strongly motivate us to conduct further research on this topic. First, in our experience, the data generation approaches studied in those three papers seem to require quite some customization for each input dataset, which is why their ease of use in daily work of pricing actuaries seems to be limited. Second, the benefit of augmenting original data with synthetic data is not addressed in those papers.
@@
Literature overview. Kuo (2020) pioneered the use of conditional tabular GANs for generating actuarial datasets, focusing on ratemaking and lapse modeling. In his experiments with the freMTPL2freq dataset, the researcher showed that generalized linear models (GLMs) trained on synthetic data achieved performance comparable to GLMs trained on real data, and the distributions of both continuous and categorical variables in the synthetic datasets closely mirrored the real datasets. However, the relative effects (how much each level of a categorical variable affects the predicted outcome relative to a reference level) of GLM models constructed using synthetic datasets do not show the same patterns as those of models constructed using real datasets. Because of the observed instability in the parameters of a GLM trained on synthetic data, developing a rating plan based on it might not be the best approach.
@@
Inspired by Kuo (2020), Côté et al. (2025) investigate three GAN-based models for generating synthetic insurance data: Multi-Categorical Wasserstein GAN with Gradient Penalty (MC-WGAN-GP), Conditional Tabular GAN (CTGAN), and Mixed Numerical and Categorical Differentially Private GAN (MNCDP-GAN). In a case study using the freMTPLfreq data, they find that MC-WGAN-GP performs best in replicating univariate distributions and the observed claim frequency for each category of categorical variables, followed by CTGAN, and then MNCDP-GAN. In assessing the consistency of GLMs fitted on real versus synthetic data, CTGAN-based GLMs most closely resemble the real-data-based GLM in terms of average estimated GLM coefficients. However, MC-WGAN-GP based GLM predictions are closest to the real-data-based GLM predictions when using Mean Absolute Error and Mean Squared Error as evaluation metrics. The authors conclude that neither of the tested models performs well across all key performance indicators (KPIs) with a high level of training data confidentiality.
@@
As an alternative to GAN-based models, Jamotton and Hainaut (2024) propose a Variational Autoencoder (VAE) model for the generation of insurance data. Their approach features a quantile transformation for continuous variables to address multimodality. In their VAE, the researchers use a reconstruction loss function that consists of two parts: cross-entropy for categorical variables and MSE for continuous variables. In a case study using a dataset from the homepage of the book Ohlsson and Johansson (2010), the researchers find that the marginal distributions of the real datasets and those of the synthetic dataset are almost the same, as supported by visual inspection, Kolmogorov-Smirnov tests for continuous variables, and Chi-Squared tests for categorical variables. Based on Pearson and Spearman correlations for continuous variables and contingency scores based on contingency tables for categorical variables, the authors conclude that synthetic data preserves pairwise relationships among variables reasonably well. Training a GLM on synthetic data generated by their VAE and noticing that the respective Poisson deviance is almost the same as the one for a GLM trained on the original data, the scientists conclude that synthetic data can substitute real data in actuarial analyses.
Fundamentally different data generation methods are those based on multivariate imputation of data. While less explored in actuarial science, these methods are well-established in statistical disclosure control. The conceptual foundation was laid by Rubin (1993), who proposed treating the creation of synthetic data as a missing data problem: synthetic values are drawn from the posterior predictive distribution of the original data. To obtain the predictive distribution of the original data, it is common to use one of the two approaches: Sequential Regression (SR) or Multivariate Imputation by Chained Equations (MICE).
The SR approach approximates the joint multivariate distribution of data by explicitly factorizing it into a strictly ordered chain of univariate conditional densities (i.e., 1st variable, then 2nd variable given the 1st variable, and so on). Reiter (2005) demonstrates that classification and regression trees (CART) can be effectively integrated into the SR framework. Building on that paper, Caiola and Reiter (2010) show that using Random Forests (RFs) as imputation models has multiple benefits for creating synthetic datasets using SR: the synthesizer requires minimal fine-tuning, naturally handles variables of any type, scales well to high-dimensional data, and automatically captures non-linear relationships. This synthetic data generation methodology is implemented in the R-package synthpop, see Nowok et al. (2016). SR approach reduces the synthesis task to a series of unidirectional regression steps according to a user-defined “synthesis order” (visit.sequence parameter in synthpop package), e.g., first Age, then Income given Age, then Family Status given Income and Age, and so on). As mentioned in Caiola and Reiter (2010) and Drechsler and Reiter (2011), there is no mathematical theory underpinning the ordering of variables for generation, so different choices of the sequence could produce synthetic datasets of different quality. Once the user fixes the synthesis sequence, earlier variables can influence the generation of the later ones, but not the other way around, which is why the dependence structure among synthetically generated variables may not be close enough to the true dependence structure and may lead to biased estimation of multivariate relationships (Reiter, 2005). Since ratemaking data can contain variables with endogenous, bidirectional dependencies (e.g., the interplay between coverage limits, deductibles, and vehicle value) that lack a natural synthesis order, we do not explore this method in our paper and focus on MICE, which is independent of the order in which variables are generated.
Developed at the end of the nineties, MICE operates by modeling the conditional density of each variable given all other variables, which is why it is also called the Fully Conditional Specification approach. This is an important difference to the SR methodology in which every single variable is imputed/generated conditional on a model trained on only the variables located before the current variable in the visiting sequence. Similarly to SR, MICE handles diverse variable types without explicitly defining a high-dimensional joint distribution, c.f., Drechsler (2011), but it offers more flexibility and is easier to use in practice because it does not require any ordering of variables for data generation. Instead, MICE utilizes an iterative Gibbs sampling scheme. By cycling through the covariates multiple times, the algorithm relies on the asymptotic convergence of the Markov Chain to reach a stationary distribution that reflects the complex, multidimensional dependence structure of the data. This approach to data imputation is implemented in the R-package mice, see van Buuren and Groothuis-Oudshoorn (2011). As with SR, MICE can also be used for synthetic data generation, which is demonstrated by Volker and Vink (2021) for a medicine-related dataset with $9$ variables and about $750$ observations. The researchers highlighted that MICE with CART as the imputation model produces indistinguishable synthetic records that support unbiased, valid statistical inferences, and offer a user-friendly solution for securely sharing and analyzing sensitive data. Although SR and MICE methodologies for synthetic data generation have been gaining popularity and credibility in the medical and social sciences, they seem to have been largely overlooked by the actuarial community in favor of deep generative models described at the beginning of our literature overview.
@@
A related idea of generating artificial data using imputation appeared in Neves et al. (2022), where the authors propose a so-called tabulator framework. In their approach, they propose to randomly (not on variable-by-variable basis) select data entries to be set to missing and then be imputed by a so-called back-end engine for imputation. As a model driving this imputation engine, the authors propose to use one of the three methods that can be seen as modifications of Generative Adversarial Imputation Networks (GAINs, originally proposed by Yoon et al. (2018)): Slim GAIN (SGAIN), Wasserstein Slim GAIN with Clipping Penalty (WSGAIN-CP), Wasserstein Slim GAIN with Gradient Penalty (WSGAIN-GP). According to the researchers, tabulator has a comparable performance (tested on 10 datasets from various fields like healthcare, economics, and computer science) with CTGAN that, in turn, is better than medical Generative Adversarial Network (medGAN), Variational Encoder Enhanced Generative Adversarial Network (VEEGAN), and Table Generative Adversarial Network (table-GAN). The researchers conclude that the tabulator preserves the dependence structure of the original dataset better than CTGAN, but both approaches struggle when generating synthetic data from original imbalanced data. It is also stated that synthetic data from tabulator improves the performance of regression/classification tasks only marginally.
Recent comprehensive reviews of literature on synthetic data generation indicate that deep generative models do not consistently outperform established methods for tabular data. Reiter (2023) observes that statistical agencies, such as the U.S. Census Bureau, continue to rely on sequential conditional modeling for major data products, citing CART as a “primary engine” or “leading contender” for synthesis tasks. Furthermore, Drechsler and Haensch (2024) report that despite the great success of GANs in computer science, comparative evaluations reveal that sequential regression based on CART frequently offers the highest utility but also the highest disclosure risk.
@@
Readers interested in the literature on insurance data generation beyond traditional ratemaking are referred to So et al. (2021) and So and Jeong (2022) for telematics data, to Gabrielli and Wüthrich (2018), Avanzi et al. (2021), Avanzi et al. (2023), Bauwelinckx et al. (2024) for reserving data, to Campo and Antonio (2023) for auto-insurance fraud-detection data, and to Chapter 10 in Wuthrich et al. (2025) for a brief introduction to generative modeling tools for actuaries. We refer readers to Fonseca and Bacao (2023) and Brophy et al. (2023) for a systematic review of the literature on the GAN-based generation of tabular data and on the GAN-based generation of time series data, respectively. For an overview comparing deep generative models, including but not limited to VAEs and GANs, we direct readers to Bond-Taylor et al. (2022).
Our contributions. First, we benchmark the utility and practicality of several amputation-imputation-based approaches to data generation for actuarial ratemaking. These techniques are based on Multivariate Imputation by Chained-Equations (MICE) with Random Forests (RFs) as imputation models. According to our tests, these approaches can be more suitable for “out-of-the-box” usage in practice than the generative models studied in Kuo (2020), Côté et al. (2025), Jamotton and Hainaut (2024), while having a comparable performance.
Second, using a publicly available French Motor Third-Part Liability dataset called freMTPL2freq (Dutang and Charpentier, 2019) as a training dataset, we compare $10$ approaches to ratemaking data generation, including the existing approaches (Kuo, 2020; Côté et al., 2025; Jamotton and Hainaut, 2024), their modifications that address the challenge of GANs in dealing with high-cardinality categorical variables, and the newly introduced amputation-imputation-based methods. We analyze the quality of synthetic datasets by studying two aspects: the similarity between the distribution of the original data and of synthetic data, the ability of methods to capture structural dependence among covariates and the response variable. Furthermore, we assess the impact of generically augmenting original data with synthetic data on the performance of GLMs for predicting claim frequency.
Our main results are threefold. First, we illustrate that the MICE-RF approach is a competitive method for generation of synthetic ratemaking data in comparison to deep generative models such as a VAE and a CTGAN. The important practical advantage of the MICE-RF algorithm is its ease of use, given the existing packages in R and the versatility of RFs in dealing with variables of different types without much pre-processing. Second, we show that modifying a CTGAN with an AE for categorical-variable pre-processing improves the quality of synthetic high-cardinality categorical variables but deteriorates the performance of this hybrid approach in terms of other criteria. In general, MICE-RF-based approaches are the best-performing in terms of most performance metrics we consider in our experiments. Third, our case studies indicate that for our experimental datasets based on the freMTPL2freq, data augmentation does not improve the performance of GLMs trained on the original data merged with synthetic one.
@@ 2 General setup
We denote a tabular dataset by $\mathcal{S}:=\left\{\bm{s}_{i}\right\}_{i=1,\dots,n_{ROD}}$, where $n_{ROD}$ is the number of rows in the original data (number of observations), $\bm{s}_{i}=(\bm{x}_{i},v_{i},y_{i})$ is a tuple with three elements such that $\bm{x}_{i}=(x_{i;1},\dots,x_{i;j},\dots,x_{i;n_{C}})^{\top}$ is a $n_{C}$-dimensional vector of observed values of covariates (independent variables), $n_{C}\in\mathbb{N}$, is the transpose operator, $v_{i}\in(0,1]$ is an exposure-to-risk (proportion of a year during which the observation $i$ was tracked) and $y_{i}$ is an observed value of a one-dimensional response variable (dependent variable) for an observation $i=1,\dots,n_{ROD}$. A tuple $\bm{s}_{i}$ can be interpreted as an insurance policy snippet, $\bm{x}_{i}$ – as a vector of policy characteristics, $v_{i}$ – as the fraction of a year during which a policy snippet $i$ was active, and $y_{i}$ – as the observed number of claims (or the monetary loss associated with a claim in the case of severity modeling). For simplicity, we may omit the index $i$ indicating the ordinal number of the observation, when it is not essential for interpretation, and just use $(\bm{x},v,y)$. The vector $\bm{x}$ can have both categorical and numerical components. We denote the number of categorical covariates in $\bm{x}$ by $n_{CC}$ and the number of numerical covariates in $\bm{x}$ by $n_{NC}$: $n_{NC}=n_{C}-n_{CC}$.
@@
In ratemaking, it is common to model the link between $y$ and $(\bm{x},v)$ via a generalized linear model (GLM). GLMs were introduced in Nelder and Wedderburn (1972) and became very popular in insurance, where they are routinely used in various business lines such as claim management, underwriting, etc. In our paper, we will only discuss the basics of a Poisson GLM, which is frequently used to predict insurance claim counts and which we use in numerical studies. For a deep dive into the theory and practice of GLMs in insurance readers are referred to Chapter 4 of Denuit and Trufin (2019) and Chapter 5 of Wüthrich and Merz (2023).
@@
In (2.3), $x_{j}$ (and/or $x_{k}$) can be just a feature representing one category. To consider an interaction with an entire categorical variable, a sum of different (2.3) (one for each category in the respective categorical variable, except for the dummy one) must be added to the right-hand side of (2.2). For more information on modeling interactions in GLMs, see, e.g., Fox and Weisberg (2018), or Section 4.2.5 in Denuit and Trufin (2019).
Let us denote the true value of $\bm{\beta}$ by $\bm{\beta}^{\ast}$. In practice, $\bm{\beta}^{\ast}$ is not known and must be estimated from collected data. Using a training data $\mathcal{S}_{train}$, an actuary can obtain an estimate of $\bm{\beta}^{\ast}$ by a so-called maximum likelihood estimation (MLE), which is equivalent to minimizing Poisson deviance in the case of a Poisson GLM. When comparing different GLM models, the one with a lower Poisson deviance on the respective test set $\mathcal{S}_{test}:=\mathcal{S}\setminus\mathcal{S}_{train}$ is considered a better-performing one. An alternative performance measure is the Akaike Information Criterion (AIC), which takes into account the number of parameters of a GLM and is based only on $\mathcal{S}_{train}$.
@@
To address the aspect of data quantity and quality, an actuary can generate a synthetic dataset $\mathcal{S}^{A}$ of a desirable size and reasonable quality. In this situation, two natural questions arise: one is related to the usage of synthetic data as a standalone set, and the other is related to working with the original dataset that is augmented with a part of the synthetic dataset. First, how well does $\mathcal{S}^{A}$ preserve (from training data $\mathcal{S}_{train}$) the marginal distributions of variables as well as the multivariate relationships among covariates, and how consistent is a GLM trained on $\mathcal{S}^{A}$ with a GLM fitted on $\mathcal{S}_{train}$? Second, which advantages and disadvantages does training GLMs on an augmented dataset $\mathcal{S}_{train}\cup\mathcal{S}^{A}$ have in terms of GLM’s predictive performance and the estimates of $\bm{\beta}^{\ast}$ (e.g., $\widehat{\bm{\beta}}(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ and/or $\widehat{\widehat{\bm{\beta}}\,}\!(\mathcal{S}_{train}\cup\mathcal{S}^{A})$ are closer to the true $\bm{\beta}^{\ast}$)? These are the two main research questions that we aim to answer in this paper.